# A Simple and Efficient Joint Measurement Strategy for Estimating Fermionic Observables and Hamiltonians

### v1

@@ VII Numerical benchmarks
For the calculations in Table 1, the coefficients $\alpha_{A}^{(r)}$ of the estimator (19) are chosen as optimal (i.e., minimizing the variance) for the ground state of a given Hamiltonian. However state-independent coefficients, uniform over rounds, can also be chosen. For details of the optimization and coefficient choice see Appendix F. Note that our estimator requires four distinct measurement rounds, and a fair comparison with strategies using a single round requires multiplying our variance by four. For further technical details see Appendix E.
@@
[…]
@@
By taking advantage of the fact that FLO circuits can be efficiently classically simulated [36, 49], we also implemented a simulation of our strategy on Gaussian states, whose energies can be easily calculated analytically (see e.g. [50]). We used the package FLOYao.jl [51] to classically simulate individual runs of our protocol, verifying (up to statistical precision) that the strategy correctly estimates the above molecular Hamiltonians on random Gaussian states. The empirical variance of this estimation also confirms (within statistical precision) the accuracy of the variance formula derived in Appendix E.
VIII Conclusions
@@
We have introduced a scheme to estimate non-commuting fermionic observables based on the implementation of a joint measurement of noisy versions of products of Majorana fermion observables. We focused on products of degree two and four, motivated by Hamiltonians from quantum chemistry. For these observables and standard benchmark Hamiltonians our scheme exhibits sample complexity guarantees comparable with fermionic classical shadows.
@@
The two distinctive features of our strategy, which make it attractive for practical purposes, are the low number of measurement settings required and the low circuit depth in certain lattices. First, in our scheme to measure all degree-2 and all degree-4 Majorana monomials with visibility $\Theta(N^{-{1/2}})$ and $\Theta(N^{-{1}})$ respectively we need at most 9 measurement settings. For the special case of electronic structure Hamiltonians relevant for quantum chemistry, as defined in Section VI, we need 4 measurement settings to estimate the energy. Second, if the system gets encoded into a 2D rectangular lattice via the Jordan-Wigner encoding, our strategy can be implemented in gate depth $\mathcal{O}(N^{1/2})$ with $\mathcal{O}(N^{3/2})$ 2-qubit gates. This compares favorably with fermionic and matchate classical shadows, which both require depth $\mathcal{O}(N)$ with $\mathcal{O}(N^{2})$ 2-qubit gates.
@@
There are several potential directions in which this work can be developed further. One interesting question to consider is how the choice of unitaries in the measurement process can be optimized for a particular Hamiltonian, in order to estimate the relevant product with higher visibility. The choice of unitary can significantly affect the variance, as demonstrated in Table 2 of Appendix F. Optimization methods have already led to substantial reductions in the sample complexity of classical shadow protocols, especially in the locally biased regime (see Table 1). Lastly, in this work we only focus on degree two and four Majorana operators. A natural extension would be to consider joint measurement strategies for higher degree Majorana observables, as well as to investigate whether other layouts, encodings, or more general orthogonal transformations provide performance benefits.
@@ Acknowledgements
The authors acknowledge financial support from the TEAM-NET project co-financed by the EU within the Smart Growth Operational Program (contract no. POIR.04.04.00-00-17C1/18-00). D.M. acknowledges support from PNRR MUR Project No. PE0000023-NQSTI. We would also like to thank Ingo Roth, Martin Kliesch and Alexander Gresch for helpful discussions, as well as Katarzyna Kowalczyk-Murynka for assistance with numerical computations.
Appendix A Variance of the Hamiltonian estimator
where $\nu^{(r)}_{A}=\det(R^{(r)}_{A,f(A)})$, $\nu^{(r)}_{A^{\prime}}=\det(R^{(r)}_{A^{\prime},f(A^{\prime})})$ and $\nu^{(r)}_{A\triangle A^{\prime}}=\det(R^{(r)}_{A\vartriangle A^{\prime},f(A)\vartriangle f(A^{\prime})})$. The matrix $R^{(r)}\in O(2N)$ corresponds to the FLO unitary $U^{(r)}$ implemented in the $r$–th round and $A\vartriangle A^{\prime}$ denotes the symmetric difference of $A$ and $A^{\prime}$, i.e., $A\vartriangle A^{\prime}=(A\cup A^{\prime})\setminus(A\cap A^{\prime})$. We have introduced the notation $f(A)\subset[2N]$ to label the subset of Majorana pairs that are measured in step (iii) of the general joint measurement protocol. For each $A$, the set $f(A)$ implicitly depends on $r$, and is chosen uniquely (for a given $r$) to optimize $\nu_{A}^{(r)}$, as described in Sec. VI.
Appendix B Number of measurement settings ensuring all quadruples are measured in 2D setting
Below we will show that at most seven ($7$) vertical permutations of the rectangular layout of modes considered in the main text suffice to ensure measurability of all Majorana quadruples (for the case when $L+1$ is a prime number). Let $\mathcal{P}$ be an arrangement of $\tilde{N}=l\cdot k$ different symbols placed in square boxes arranged in an $l\times k$ rectangular pattern (note that in what follows $l$ plays the role of $L+1$ from the main text). Let
@@ Appendix B Number of measurement settings ensuring all quadruples are measured in 2D setting
Appendix C Shallow circuits for vertical FLO unitaries in 2D layouts under the JW encoding
In this appendix we prove that structured FLO transformations $U_{\mathrm{pair}}$ and $U_{\mathrm{resh}}^{(r)}$ introduced in Section IV to estimate Majorana pairs and quadruples can be implemented in depth $\mathcal{O}(N^{1/2})$ and using $\mathcal{O}(N^{3/2})$ nearest neighbour two-qubit gates arranged on a rectangular lattice.
In the following we will be using a "snake" labelling, used e.g. in [44, 47], of qubits that encode fermions via the Jordan-Wigner transformation. In this encoding a rectangular $l_{x}\times l_{y}$ grid of qubits encodes $2N=2l_{x}l_{y}$ Majorana modes (see Fig. 5a for an example of an arrangement of qubits and labelling of modes for $l_{x}=2,l_{y}=5$). By horizontal encoding we mean that qubits are arranged along a "snake" which traverses subsequent rows in the rectangular grid horizontally (see Fig. 5a). By vertical encoding we mean that qubits are arranged along a "snake" which traverses subsequent columns vertically (see Fig. 5b).
Figure 5: Graphical presentations of Majorana labellings and vertical or horizontal qubit encodings used in this Appendix, as well as a visual presentation of vertical FLO transformations (part b)). Parts a) and c) present different ways of labelling $2N=2l_{x}l_{y}$ Majorana modes in $l_{x}\times l_{y}$ rectangular qubit lattice for $l_{x}=2,l_{y}=5$. Dots in the figures represent Majorana modes while qubits are represented by blue rounded rectangles. The green "snakes" in a) and b) represent a horizontal and vertical encoding respectively. Part a) presents a snake-like labelling used in [44, 47] - it is compatible with the horizontal Jordan-Wigner encoding. Part c) presents the labelling used in the main part of this work. Finally, part b) presents the locality structure (indicated by two green snakes traversing through columns of modes 1-2 and 3-4 respectively) needed to implement a vertical FLO transformation $V=V_{I}V_{II}$. Furthermore, the combined path (of the "snake" traversing both qubit columns) realizes the vertical Jordan-Wigner encoding required in the proof of Proposition 3.
Remark. The "snake" labelling of Majorana modes differs from the one considered in the main part of this work (cf. Fig. 5c). However, both labellings can be related by a horizontal transformation of the modes $P=\oplus_{\alpha-\text{even}}O_{\alpha}$, which reverses the order of Majorana modes in even rows (for the example with $l_{x}=2,l_{y}=5$, we have that $O_{2}$ realizes $\sigma_{2}=(5,8)(6,7)$ and $O_{4}$ realizes $\sigma_{4}=(13,16)(14,15)$, where we used (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) to denote a transposition. Since $P$ does not mix Majorana operators in different rows, the corresponding unitary transformation $V_{P}$ can be implemented in depth $d_{p}=\mathcal{O}(l_{x})$ and using $n_{P}=\mathcal{O}(l_{x}l_{y}^{2})$ nearest neighbour two qubit gates. Consequently, the (asymptotic) bounds on the depth and gate-count for implementing $U_{\mathrm{pair}}$ and $U_{\mathrm{resh}}^{(r)}$ are identical in both labellings of qubits.
Definition.
(Vertical FLO transformations) Consider a snake-like labelling of $2N=2l_{x}l_{y}$ Majorana modes in an $l_{x}\times l_{y}$ array of qubits (cf. Fig. 5a). Let $C_{i}$ ($i\in[2l_{x}]$) denote the subset of labels of Majorana modes belonging to the i’th column of the rectangular arrangement. Furthermore, let $\mathcal{P}_{\beta}=C_{2\beta-1}\cup C_{2\beta}$ ($\beta\in[l_{x}]$) and $W_{i}=\mathrm{span}_{\mathbb{R}}\{\{\left|j\right>\ |\ j\in\mathcal{P}_{i}\}\}$. We say that $V\in\mathrm{FLO}(N)$ is vertical if the corresponding orthogonal transformation has the form
$O=\bigoplus_{\beta=1}^{l_{x}}O_{\beta},$ (28)
See Fig. 5b for a depiction of the vertical FLO transformation for $l_{x}=2,l_{y}=5$.
The compilation procedure presented in the proof below can likely be made more efficient by taking into account the structure of the unitaries $U_{\mathrm{pair}}$, $U_{\mathrm{resh}}^{(r)}$, as well as a possible further improvement in the compilation of general vertical FLO transformations.
Proposition 3.
Let $V\in\mathrm{FLO}(N)$ be a vertical FLO transformation on a $l_{x}\times l_{y}$ array of qubits. Then $V$ can be implemented in depth $d_{V}=\mathcal{O}(l_{x})+\mathcal{O}(l_{y})$ and using $n_{V}=\mathcal{O}(l_{y}^{2}l_{x})$ nearest neighbour two qubit gates, assuming one extra layer of qubits is available per row.
The above proposition shows the desired depths and gate-counts necessary to implement $U_{\mathrm{pair}}$ and $U_{\mathrm{resh}}^{(r)}$ in a rectangular array of qubits. Indeed both unitaries are vertical FLO transformations and for $l_{x}=\mathcal{O}(N^{1/2})$, $l_{y}=\mathcal{O}(N^{1/2})$ we have gate depth $d_{V}=\mathcal{O}(N^{1/2})$ and nearest neighbour gate-count $n_{V}=\mathcal{O}(N^{3/2})$.
Proof of Proposition 3.
We will follow the proof strategy presented in [44] for the case of vertical passive FLO transformations. The idea is to map, via a suitable unitary transformation $\Gamma$, between horizontal and vertical realizations of the Jordan-Wigner encoding – see Fig. 5 a and b for an example for $l_{x}=2,l_{y}=5$.
We first note that in the vertical encoding, an implementation of an arbitrary vertical $V$ can be achieved in $\mathcal{O}(l_{x})$ depth and by using $\mathcal{O}(l_{x}^{2}l_{y})$ gates. This is because each $O_{\beta}$ in the decomposition (28) can be realized by a product of $\mathcal{O}(l_{y}^{2})$nearest neighbour" rotations, $\mathrm{exp}(i\varphi\gamma_{a}\gamma_{b}$), where $a,b$ are indices of neighbouring Majorana modes (see e.g. [44] or [45] for the detailed proof). In the vertical encoding, such transformations are realized by single qubit Pauli $Z$ rotations, $\mathrm{exp}(i\varphi Z_{i})$, $i\in[N]$ (they realize transformations $\mathrm{exp}(i\varphi\gamma_{2i-1}\gamma_{2i})$) and local gates generated by the following local couplings
where $\alpha,\beta$ are neighbouring qubits that belong to the same column in the qubit arrangement. In the horizontal encoding, single qubit Pauli rotations still generate $\mathrm{exp}(i\varphi\gamma_{2i-1}\gamma_{2i})$, while generators from $\mathcal{C}_{\alpha\beta}$ have to be "dressed" by an additional parity term $\prod_{\alpha<i<\beta}Z_{i}$ in order to generate the right fermionic transformation. The above discussion shows that if we can find $\Gamma$ such that for all $i\in[N]$ we have $\Gamma Z_{i}\Gamma^{\dagger}=Z_{i}$ and, furthermore, for all neighbouring qubits $\alpha,\beta$ that belong to the same column, and for all $A\in\mathcal{C}_{\alpha\beta}$, we have
$\Gamma A\Gamma^{\dagger}=A\prod_{\alpha<i<\beta}Z_{i}\ ,$ (30)
then we have
$\Gamma V_{v}\Gamma^{\dagger}=V\ ,$ (31)
where $V_{v}$ is the circuit realizing an FLO transformation which in the horizontal snake-like encoding realized by $V$. In what follows we explicitly construct a transformation $\Gamma$ that satisfies the above conditions and can be realized in depth $d_{\Gamma}=\mathcal{O}(l_{x})+\mathcal{O}(l_{y})$, by using $n_{\Gamma}=\mathcal{O}(l_{x}l_{y})$ nearest neighbour two-qubit gates, and a single auxiliary qubit per row. This, together with condition (31), gives the claimed estimates for the depth and gate counts required to implement $V$:
Figure 6: A graphical presentation of a circuit implementing a unitary $\Gamma$ for $l_{x}=3$, $l_{y}=3$. Black and blue dots represent the system and ancilla qubits respectively. The subsequent diagrams depict subsequent stages of the computation. A detailed explanation of different parts of the circuit is given in the last part of the proof of Proposition 3. Specific steps in the figure are referenced in text by their row (I-V) and column (A-D).
Before presenting a circuit that implements the desired unitary $\Gamma$, we first make some general observations which will be useful in later analysis. We first note that if $\Gamma$ is diagonal in the computational basis, then $\Gamma Z_{i}\Gamma^{\dagger}=Z_{i}$ and it suffices to check the condition from Eq. (30). This condition is equivalent to requiring that for all neighbouring qubits $\alpha,\beta$ belonging to the same column, for all $A\in\mathcal{C}_{\alpha\beta}$, and for arbitrary basis vectors $\left|s\right>,\left|s^{\prime}\right>$ such that $s\right|A\left|s^{\prime}\right>\neq 0$, we have
where $s=s_{1}s_{2}\ldots s_{N}\ ,\ s^{\prime}=s^{\prime}_{1}s^{\prime}_{2}\ldots s^{\prime}_{N}$ are bitstrings corresponding to $\left|s\right>,\left|s^{\prime}\right>$ and $\lambda_{s},\lambda_{s^{\prime}}$ are the corresponding eigenvalues of $\Gamma$. For every $\mathcal{C}_{\alpha\beta}$, the characterization of bitstrings $s,s^{\prime}$ such that $s\right|A\left|s^{\prime}\right>\neq 0$ is straightforward since $A\in\mathcal{C}_{\alpha\beta}$ act as simultaneous bitflips on the computational basis states (up to an overall phase). Consequently we get that bitstrings $s,s^{\prime}$ for which (34) has to be satisfied are of the following form (for any pair of qubits $\alpha,\beta$ which belong to the same column and are neighbours)
[…]
The above procedure can be parallelized by operating simultaneously on disjoint pairs of rows in both stages of the computation. Furthermore, the same procedure can be easily generalized to arbitrary $l_{x}$ and $l_{y}$. After a simple count, we obtain the claimed bounds on the depth and two-qubit gate counts, namely $d_{\Gamma}=\mathcal{O}(l_{x})+\mathcal{O}(l_{y})$ and $n_{\Gamma}=\mathcal{O}(l_{y}l_{x})$.
Appendix D Four measurement rounds are sufficient to estimate Hamiltonians

### v2

@@ II.5 Numerical benchmarks
Table 1: Comparison of variances for estimating energies of standard benchmark molecules. The variances are calculated for the ground states of five molecular Hamiltonians, with units Ha2. The first four strategies (taken from a summary in [1]) include: classical shadows with Pauli measurements (CS-Pauli [24]); locally biased classical shadows (LBCS [25]); basis-rotation grouping (BRG [35]); and classical shadows with fermionic Gaussian unitaries (CS-FGU [1]). Where applicable ([24, 25, 35]), the Jordan-Wigner encoding was used. The last column summarises our results, optimized over the coefficients $\alpha_{A}^{(r)}$.
For the calculations in Table 1, the coefficients $\alpha_{A}^{(r)}$ of the estimator (14) are chosen as optimal (i.e., minimizing the variance) for the ground state of a given Hamiltonian. However state-independent coefficients, uniform over rounds, can also be chosen. For details of the optimization and coefficient choice see Appendix E. Note that our estimator requires four distinct measurement rounds, and a fair comparison with strategies using a single round requires multiplying our variance by four. For further technical details see Appendix F.
@@
By taking advantage of the fact that FLO circuits can be efficiently classically simulated [38, 45], we also implemented a simulation of our strategy on Gaussian states, whose energies can be easily calculated analytically (see e.g. [46]). We used the package FLOYao.jl [47] to classically simulate individual runs of our protocol, verifying (up to statistical precision) that the strategy correctly estimates the above molecular Hamiltonians on random Gaussian states. The empirical variance of this estimation also confirms (within statistical precision) the accuracy of the variance formula derived in Appendix D.
III Discussion
@@
We have introduced a scheme to estimate non-commuting fermionic observables based on the implementation of a joint measurement of noisy versions of products of Majorana fermion observables. We focused on products of degree two and four, motivated by Hamiltonians from quantum chemistry. For these observables and standard benchmark Hamiltonians, our scheme exhibits sample complexity guarantees comparable with fermionic classical shadows.
@@
The three distinctive features of our strategy, which make it attractive for practical purposes, are the low number of complicated measurement settings required, low circuit depth in certain lattices, and the dependence of the observables’ estimates on single-qubit measurements on few qubits. First, in our scheme to measure all degree two and four Majorana monomials with visibility $\Theta(N^{-{1/2}})$ and $\Theta(N^{-{1}})$, respectively, we need at most nine measurement settings. For the special case of electronic structure Hamiltonians relevant for quantum chemistry, as defined in Sec. II.4, we need four measurement settings to estimate the energy. In both cases an additional randomization over Majorana products is required, however this can be implemented easily, by a single gate layer under any fermion-to-qubit mapping. Second, if the system gets encoded in a 2D rectangular lattice via the Jordan-Wigner encoding, our strategy can be implemented in gate depth $\mathcal{O}(N^{1/2})$ with $\mathcal{O}(N^{3/2})$ 2-qubit gates. This compares favorably with fermionic and matchgate classical shadows, which both require depth $\mathcal{O}(N)$ with $\mathcal{O}(N^{2})$ 2-qubit gates. Third, our strategy uses single-qubit measurement outcomes from only one or two qubits to estimate Majorana pairs or quadruples, respectively. As such it may readily be combined with randomized error mitigation techniques [48].
@@
There are several potential directions in which this work can be developed further. One interesting question to consider is how the choice of unitaries in the measurement process, discussed in Sec. II.3, can be optimized for a particular Hamiltonian, in order to estimate the relevant products with higher visibility. The choice of unitary can significantly affect the variance, as demonstrated in Table 2 of Appendix E. Optimization methods have already led to substantial reductions in the sample complexity of classical shadow protocols, especially in the locally biased regime (see Table 1). Moreover, in this work we only focus on degree two and four Majorana operators. A natural extension would be to consider joint measurement strategies for higher degree Majorana observables, as well as to investigate whether other layouts, fermion-to-qubit encodings, or more general orthogonal transformations, provide performance benefits. Furthermore, the effect of noisy physical implementations on the scheme could be studied, similarly to [30]. Finally, our strategy can be directly applied on native fermionic computing or simulation platforms, e.g. programmable neutral atom arrays [49], without the need to map to qubits, and is naturally compatible with devices using a 2D layout. An interesting research direction would be to experimentally apply it on such devices.
IV Materials and methods
IV.1 Proof for Joint Measurement strategy
Below we provide proof of Proposition 1:
Proof.
We first note that
where we have used (3) and the commutation relation of the operators $\gamma_{X}$ and $\gamma_{Y}$. In order to prove the claim in the proposition it suffices to show that for $x=\pm 1$ we have $\mathrm{Pr}_{X,\bm{q}}(e_{A}=x)=\operatorname{tr}(\rho\frac{1}{2}(\mathbb{I}+x\cdot\eta_{A}\gamma_{A}))$, where $X$ and $\bm{q}$ are random variables on which $e_{A}$ depends. Using the definition of $e_{A}$ in step (iv) and the fact that the distribution of $X$ is uniform we have
$\frac{s_{AB}}{2^{2N+1}}\sum_{\begin{subarray}{c}X,Y\subseteq[2N]\\ |Y|=|B|\end{subarray}}\operatorname{tr}\left(\rho(-1)^{|A\cap X|+|X\cap Y|}\gamma_{Y}\det(R_{Y,B})\right)$
$=\frac{1}{2}\eta_{A}\operatorname{tr}(\rho\gamma_{A})\ ,$ (19)
where we have used that for $|A|,|B|$–even:
IV.2 Implementation of measurement strategies
Figure 4: Two exemplary realizations of the Jordan-Wigner encoding in systems of qubits that exhibit different spatial locality: a) 1D array and b) $L/2\times(L+1)$ rectangular qubit array (for $L=4$). The individual qubits are indicated by consecutive pairs within the blue rounded rectangles labelled by $Q_{1},Q_{2},\ldots$, each enclosing two Majorana mode operators indicated by black dots. Importantly, locality in Majorana modes (e.g. coupling $\mathrm{exp}(\mathrm{i}\varphi\gamma_{i,j})$ for neighboring Majorana modes $i,j$) does not in general correspond to locality with respect to the structure of the qubit lattice.
In this part we discuss the gate depths and numbers required to implement our joint measurement strategy on a quantum computer in different fermion-to-qubit mappings and qubit layouts.
General strategy. In standard fermion-to-qubit mappings (like Jordan-Wigner [41], Parity [50], Bravyi-Kitaev [39], or superfast encodings [39, 51]; see [50] for a review), unitaries from step (i) of Sec. II.2 can be implemented via a depth-1 circuit consisting of Pauli gates. On the other hand, the realization of any $U\in\mathrm{FLO}(N)$, including those from step (ii), can be achieved with a circuit of depth at most $d_{\mathrm{gen}}=\mathcal{O}(N^{3})$ and $n_{\mathrm{gen}}=\mathcal{O}(N^{4})$ two-qubit gates in any of the encodings. This follows from two observations. First, an arbitrary $U\in\mathrm{FLO}(N)$ can be realized by $U=\tilde{U}\gamma_{1}^{s}$, where $\tilde{U}$ is parity preserving (i.e., it commutes with $Q:=\gamma_{[2N]}$), and $\gamma_{1}^{s}$ ($s=0,1$) is responsible for the (optional) change of the parity. It is known that $\tilde{U}$ can be implemented using $\mathcal{O}(N^{2})$ gates of the form $\mathrm{exp}(\mathrm{i}\varphi\gamma_{i,i+1})$ [52, 53]. Second, each of the unitaries $\mathrm{exp}(\mathrm{i}\varphi\gamma_{ij})$ can be realized in depth $\mathcal{O}(N)$, which is a consequence of the fermion-to-qubit encoding that maps $\gamma_{ij}$ to a Pauli string with support on at most $\mathcal{O}(N)$ qubits, and is thus realizable in depth $\mathcal{O}(N)$ and using $\mathcal{O}(N)$ gates (unitaries $\mathrm{exp}(\mathrm{i}\varphi P)$, for a Pauli string $P$ supported on $k$ qubits can be implemented via $\mathcal{O}(k)$ two-qubit gates [54]).
However, when using the standard Jordan-Wigner encoding on a line (see Fig. 4a) it is possible to realize any $U\in\mathrm{FLO}(N)$ using $\mathcal{O}(N)$ layers of $\mathrm{exp}(\mathrm{i}\varphi\gamma_{i,i+1})$ applied in parallel and forming a brickwork pattern on a triangle layout [53]. Furthermore, in this encoding $\mathrm{exp}(\mathrm{i}\varphi\gamma_{i,i+1})$ can be realized by a constant depth unitary acting on neighboring qubits. Rotations corresponding to disjoint Majorana pair operators belonging to each layer can be realized in parallel (see [52] or [53]). This yields the overall depth $d_{\mathrm{JW}}=\mathcal{O}(N)$ and the two-qubit gate-count $n_{\mathrm{JW}}=\mathcal{O}(N^{2})$.
A rectangular qubit grid in the Jordan-Wigner encoding. Due to the particular structure we impose on the unitaries in step (ii), our joint measurement scheme for Majorana pairs and quadruples can be realized with a smaller circuit depth and using fewer two-qubit gates in a 2D layout of qubits. Consider a $L/2\times(L+1)$ lattice of qubits realizing $L(L+1)$ Majorana modes via the Jordan-Wigner encoding (see Fig. 4b). In Sec. II.3 we described specific FLO unitaries $U_{\mathrm{pair}},U_{\mathrm{sup}},U^{(r)}_{\mathrm{resh}}$ that define measurement rounds of the protocol from Prop. 2. Importantly, for the variant of the Jordan-Wigner encoding in which the Majorana modes are labelled as in Fig. 4b, $U_{\mathrm{sup}}$, $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ can all be implemented in depth $d_{2D}=\mathcal{O}(N^{1/2})$ and with $n_{2D}=\mathcal{O}(N^{3/2})$ two-qubit gates. In what follows we sketch the reasoning behind these estimates.
First, the superposition unitary $U_{\mathrm{sup}}$ is horizontal (see (12)) in the 2D Majorana layout in Fig. 2 (and consequently Fig. 4b). Such transformations are consistent with the ordering that gives rise to the JW encoding and can therefore be implemented by a (tensor) product of $L+1$ independent FLO unitaries, each acting on qubits arranged in a given row. Therefore $U_{\mathrm{sup}}$ can be implemented in depth $\mathcal{O}(N^{1/2})$ and $\mathcal{O}(N^{3/2})$ nearest-neighbor two-qubit gates.
It is more challenging to implement the unitaries $U_{\mathrm{pair}}$ and $U_{\mathrm{resh}}^{(r)}$. Note that, as explained in Sec. II.3, they are purely vertical (see (12)). That is, the permutations $\pi$ and $\sigma^{(r)}$ that correspond to $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ respectively, correspond to orthogonal transformations $R_{\pi/\sigma^{(r)}}=\oplus_{\beta=1}^{L/2}\tau^{\beta}$, where $\tau^{\beta}$ only permutes elements in the $\beta$–th column of qubits in the 2D layout under JW encoding. Importantly, in [52] (see also [55]), it was shown that such a class of FLO gates can be implemented in depth $\mathcal{O}(L)=\mathcal{O}(N^{1/2})$ and using $\mathcal{O}(L^{3})=\mathcal{O}(N^{3/2})$ two-qubit gates, by utilizing one extra auxiliary qubit per row. In Appendix G we provide the proof that these scalings hold for general FLO transformations, extending the treatment of number preserving FLO transformations in [52]). Thus, $U_{\mathrm{pair}}$ and $U^{(r)}_{\mathrm{resh}}$ can also be implemented with the claimed gate depth and count.
@@ Acknowledgements
We would like to thank Ingo Roth, Martin Kliesch and Alexander Gresch for helpful discussions, as well as Katarzyna Kowalczyk-Murynka for assistance with numerical computations. The authors acknowledge financial support from the TEAM-NET project co-financed by the EU within the Smart Growth Operational Program (contract no. POIR.04.04.00-00-17C1/18-00). D.M. acknowledges support from PNRR MUR Project No. PE0000023-NQSTI. The authors declare that they have no competing interests.
Appendix A Number of measurement settings ensuring all quadruples are measured in 2D setting
Below we will show that at most seven ($7$) vertical permutations of the rectangular layout of modes considered in the main text suffice to ensure measurability of all Majorana quadruples (for the case when $L+1$ is a prime number). Let $\mathcal{P}$ be an arrangement of $\tilde{N}=l\cdot k$ different symbols placed in square boxes arranged in an $k\times l$ rectangular pattern (note that in what follows $l$ plays the role of $L+1$ from the main text). Let
@@ Appendix A Number of measurement settings ensuring all quadruples are measured in 2D setting
Appendix B Four measurement rounds are sufficient to estimate Hamiltonians